OpenStudy (anonymous):

Given the following endpoint (A) and midpoint (M), find the other endpoint (B). M(2, 3) A(0, -1) 1) (7, 4) 2) (3, 2) 3) (-1, 0) 4) (4, 7)

OpenStudy (anonymous):

The midpoint of an interval is given by\[(a,b)=(\frac{x_1+x_2}{2})+(\frac{y_1+y_2}{2})\]Hence, we can identify,\[a=\frac{x_1+x_2}{2}\]and\[b=\frac{y_1+y_2}{2}\]In this case, we have,\[2=\frac{0+x}{2} \rightarrow x=4\]and\[3=\frac{-1+y}{2} \rightarrow y=7\]So the other point is (4,7).